1. Calculate g(b) and g′(b) where g(x)g(x) is the inverse of f(x)=3x+11.
g′(x) =

2.  Calculate g(b) and g′(b) where g(x) is the inverse of f(x)=3/(x+4), where b=3/10.
g(b)=

g'(b)=

3. Calculate g(b) and g′(b) where g(x) is the inverse of f(x)=3x+11.
g′(x) =

help please
Answer #1

Pims f(x) = 3x tll => y =33 +11 => y-leo 3 CA :: g(y) = y.!! 3 = g(x) = X-M1 3 I da lidi 3 dx Nowglx) (7) } Joy dsc lineari